a right triangle has legs lengths of 9 and 12: The altitude of the hypotenuse has lenght of ?
Answer by richwmiller(17219)

Are you looking for the length of the hypotenuse? It is 15
this is a triple multiple of the 3,4,5 right triangle
3*3=9
3*4=12
3*5=15